# AlignXIE: Improving Multilingual Information Extraction by Cross-Lingual Alignment

### v1

@@ Abstract
Empirical evidence suggests that LLMs exhibit spontaneous cross-lingual alignment. Our findings suggest that although LLMs also demonstrate promising cross-lingual alignment in Information Extraction, there remains significant imbalance across languages, revealing an underlying deficiency in the IE alignment. To address this issue, we propose AlignXIE , a powerful code-based LLM that significantly enhances cross-lingual IE alignment through two strategies. Firstly, AlignXIE formulates IE across different languages, especially non-English ones, as code generation tasks, standardizing the representation of various schemas using Python classes to ensure consistency of the same ontology in different languages and align the schema. Secondly, it incorporates an IE cross-lingual alignment phase through a translated instance prediction task proposed in this paper to align the extraction process, utilizing ParallelNER, an IE bilingual parallel dataset with 257,190 samples, generated by our proposed LLM-based automatic pipeline for IE parallel data construction, with manual annotation to ensure quality. Ultimately, we obtain AlignXIE through multilingual IE instruction tuning. Although without training in $9$ unseen languages, AlignXIE surpasses ChatGPT by 30.17% and SoTA by 20.03%, thereby demonstrating superior cross-lingual IE capabilities. Comprehensive evaluations on 63 IE benchmarks in Chinese and English under various settings, demonstrate that AlignXIE significantly enhances cross-lingual and multilingual IE through boosting the IE alignment.
@@
AlignXIE: Improving Multilingual Information Extraction by Cross-Lingual Alignment
@@ 1 Introduction
CoNLL 2003 CoNLL 2003-zh
Flan-T5 12.71 2.13
InstructUIE 92.94 20.39
Table 1: Results on CoNLL 2003 and CoNLL 2003-zh, the latter being the constructed Chinese parallel dataset.
@@
Multilingual Information Extraction (Multilingual IE) focuses on automatically extracting structured knowledge from unstructured or semi-structured texts across various languages, following manually designed schemas. Large Language Models (LLMs) (Brown et al., 2020; Touvron et al., 2023a, b), trained on massive multilingual corpora, have significantly advanced multilingual language processing. Previous studies (Zhang et al., 2023; Qi et al., 2023; Wang et al., 2024; Gao et al., 2024) have shown that LLMs exhibit spontaneous cross-lingual alignment to facilitate the transfer of abilities and knowledge across languages. Our findings suggest the presence of this alignment in IE, indicating a strong potential for improving the IE cross-lingual transfer. We first define IE parallel data, which is across various languages that share the same schema, sentences, and extracted instances. Previous studies proposed label projection to generate such data to alleviate the challenges of IE in low-resource Kolluru et al. (2022); Hennig et al. (2023); Rios et al. (2024). By translation and manual annotation, we constructed an IE parallel dataset in Chinese based on CoNLL 2003 Sang and Meulder (2003), to evaluate the InstructUIE Wang et al. (2023a). Table 1 shows the results. After training Flan-T5 Chung et al. (2022) on English IE to obtain InstructUIE, there was a notable enhancement in Chinese IE ($2.13$$\rightarrow$20.3920.3920.3920.39), strongly supporting the presence of IE cross-lingual alignment. However, the significant performance gaps between languages ($92.94$ vs $20.39$) indicate that the alignment remains weak. To address this issue, we propose two strategies to enhance IE cross-lingual alignment by aligning the schema and extraction process in different languages, thereby improving the cross-lingual and multilingual IE.
Firstly, we formulate IE tasks across different languages, particularly non-English languages, into a unified framework of code generation. Traditional multilingual IE faces a major challenge from cross-lingual schema differences, requiring schema alignment before knowledge sharing, which hinders cross-lingual transfer. Yang et al. found that unified schema representations in a single language can significantly enhance model generalization, and thus, we construct a unified schema representation to align cross-lingual schemas. Leveraging the language-agnostic nature of code, we use the Python classes to represent schemas across various languages, which ensures consistent representation of the same or similar ontology in different languages, thereby facilitating knowledge transfer for IE. The consistent schema allows the model to learn and efficiently share the knowledge of the same ontology across languages.
@@
Secondly, we incorporate an IE cross-lingual alignment instruction tuning phase. We propose a task utilizing the IE parallel data, aiming at predicting target language instances to facilitate the alignment of the extraction process. Leveraging prediction tasks with parallel data to enhance alignment through instruction tuning has become a common paradigm in multilingual processing (Zhu et al., 2023; Gao et al., 2024). Figure 1 presents an example. Similar to the label projection, the challenge of the task is to ensure that the predicted instances remain semantically consistent with the source-language ones while appearing in the target language sentence. In contrast to predicting complete IE parallel data directly, which mainly focuses on the alignment of the target-language sentence that will be provided in IE, our method emphasizes the alignment of translated instances, which is the core objective of IE cross-lingual alignment.
@@
To use high-quality IE parallel data for the phase, we propose an automatic three-stage LLM-based pipeline that implements contextual translation and multi-grained rephrasing. The pipeline achieved $99\%$ average faithfulness across $10$ languages evaluated on the label projection benchmark WikiANN Pan et al. (2017). Utilizing the pipeline, we construct a high-quality and diverse NER English-Chinese parallel dataset ParallelNER, using the GPT-4 series model supplemented with manual annotation. By distilling the IE alignment capabilities of advanced proprietary models, the IE cross-lingual alignment phase enhances cross-lingual generalization in IE tasks.
@@
Ultimately, we obtain the AlignXIE through multilingual IE instruction tuning. We evaluate the AlignXIE on $63$ benchmarks for the NER, RE, ED, and EAE tasks in Chinese and English. AlignXIE exhibits superior cross-lingual generalization across typologically diverse languages, and surpasses ChatGPT by $30.17\%$ and SoTA by $20.03\%$. In the supervised evaluation, AlignXIE consistently ranks within the top-$2$ results across $40$ (of $42$) benchmarks, with an average improvement of $3.03$ and $2.92$ points on the English RE and ED tasks, respectively. Notably, AlignXIE achieves SoTA across all Chinese IE benchmarks under all settings, effectively demonstrating the strength of knowledge transfer from English IE through our method.
Our contributions can be summarized as:
@@
  • The code-based multilingual IE method of AlignXIE unifies the representation of schemas across different languages, thereby boosting cross-lingual transfer.
@@
  • AlignXIE incorporates a cross-lingual alignment phase for IE to improve cross-lingual transfer. This phase involves instruction tuning on a newly proposed translated instance prediction task.
@@
  • We propose an LLM-based automatic IE parallel data construction pipeline. Using this pipeline, we constructed a large-scale, high-quality bilingual parallel NER dataset ParallelNER with $257,190$ samples to provide valuable resources to the community.
@@ 2 Code-Based Multilingual IE
AlignXIE formulates the IE tasks across various languages using a unified code generation framework. In this section, we introduce the instruction and completion of the proposed code-based multilingual IE. The instruction consists of two parts (§2.1): (1) representation code of schema across different languages (§2.1.1); (2) task prompts for completion (§2.1.2). Subsequently, we present the completion format (§2.2).
@@ 2.1 Instruction Format
2.1.1 Schema Representation
Class Definition
The IE schema comprises multiple concepts, each of which consists of a name and a set of attributes. The three fundamental concepts of IE schemas [Entity, Relation, Event] are first defined as base classes. For each concept of a schema, a corresponding class is subsequently defined based on its name. Each concept inherits from its respective base class by default. Then, we define the attributes of each concept (such as argument roles) as input arguments of the constructor __init__(). Specially, for IE in non-English languages, we first create a mapping from the original schema to the English schema through translation or manual annotation. The mapping standardizes the representation of the same ontology across different languages by utilizing the unified Python class. For example, the entity concept {CJK}UTF8mj사람 in Korean and {CJK}UTF8gkai人物 in Chinese are both mapped to the class PER(Entity), ensuring consistent semantic representation across linguistic boundaries to align the schema in different languages.
Class Comments
We utilize class comments to provide clear definitions of concepts similar to GoLLIE Sainz et al. (2024) and KnowCoder Li et al. (2024). Class comments comprise two components: 1) a description that explains the corresponding concept; 2) examples contain instances corresponding to the concept. For the description part, we follow the annotation guidelines of datasets by default. We propose an LLM-based approach to generate descriptions for datasets where guidelines are unavailable. To enable the LLM to better understand the concepts and the annotation style of the dataset, we first sample $10$ instances to prompt the LLM to generate an initial description. Subsequently, to avoid providing excessive examples at once to hinder effective induction, we sample $20$ instances for the LLM to iteratively refine the description based on each instance. Specifically, for each instance, if it cannot be categorized to the corresponding concept according to the description, the LLM will adjust the inappropriate parts of the current description accordingly. We use the GPT-4o-2024-08-06 as the LLM. The prompts are shown in Appendix C. For the examples part, we sample up to $10$ instances for each concept by frequency.
2.1.2 Task Prompt
The task prompt contains a docstring that defines the task and incorporates a string variable sentence in Python, which contains the text to be extracted. Then, the task prompt is concatenated to the schema, resulting in the instruction.
@@ 2.2 Completion Format
The completion represents the final output generated by the model, which begins after the assignment operation (i.e., results =). Each completion is a list, where each element is an instantiated object, representing the extracted structured knowledge instance. Additionally, we utilize the exec() function in Python to execute the instruction and completion of all code-based data in this work to ensure quality. We also use this function to execute the generated code to extract the results.
3 Training Framework
AlignXIE adopts a two-phase training framework based on instruction tuning. The first phase is the IE cross-lingual alignment phase, where we propose the task and data (§3.1). Subsequently, we conduct the multilingual IE instruction tuning to obtain AlignXIE  (§3.2). Appendix D shows examples of instruction tuning data of the two phases.
3.1 IE Cross-Lingual Alignment Phase
First, we introduce the instruction and completion of the translated spans prediction task (§3.1.1). To obtain high-quality IE parallel data for this training phase, we propose a three-stage automatic pipeline for constructing IE parallel data (§3.1.2) and create a dataset ParallelNER (§3.1.3).
3.1.1 Translated Instances Prediction Task
@@
       // Generate prompt for Contextual Translation.
@@
                   // Generate prompt for Span Rephrase. $i^{cor}_{m}=$ LLM ($prompt$) ;
@@
       $prompt=$ TemplateS ($I^{tgt},s^{src}$) ; // Generate prompt for Sentence Rephrase. $s^{tgt}=$ LLM ($prompt$); return $s^{tgt}$ ;
@@
Algorithm 1 Span Translation
3.1.2 IE Parallel Data Construction Pipeline
The construction of IE parallel data involves two critical components: span translation and schema translation. Spans refer to the desirable spans annotated in the text of the IE task Kripke and Munitz (1971); Chen and Yuille (2004). Our pipeline mainly performs span translation, followed by a post-processing step that translates the schema and associates it with the corresponding spans Onyshkevych (1994); Milward and Thomas (2000); Lu et al. (2022). Schema translation can be accomplished using LLMs and manual annotation. The primary challenge lies in the span translation, which involves two key issues: 1) Inaccurate translation: Either the spans or the sentences are inaccurately translated; 2) Missing Spans: The translated spans may be missing from the translated sentence. To address these issues, we propose a three-stage pipeline implementing contextual translation and multi-grained rephrasing, which is capable of generating parallel data for any IE task across any languages, as shown in Algorithm 1. The detailed prompts are shown in Appendix C.
Stage 1: Contextual Translation
We conduct the contextual translation first to ensure the semantic consistency of the translated sentence and spans. Given the sentence $s^{src}$ and spans list $I^{src}=\{i^{src}_{1},i^{src}_{2},...,i^{src}_{n}\}$ in the source language, the aim of this stage is to obtain the initial sentence $s^{tgt}$ and spans list $I^{tgt}=\{i^{tgt}_{1},i^{tgt}_{2},...,i^{tgt}_{n}\}$ in the target language. We use a predefined template to format the sentence $s^{src}$ with spans $I^{src}$ as the prompt for the LLM to translate jointly.
Stage 2: Span Rephrase

### v2

# KnowCoder-X: Boosting Multilingual Information Extraction via Code

@@ Abstract
Empirical evidence indicates that LLMs exhibit spontaneous cross-lingual alignment. However, although LLMs show promising cross-lingual alignment in IE, a significant imbalance across languages persists, highlighting an underlying deficiency. To address this, we propose KnowCoder-X, a powerful code LLM with advanced cross-lingual and multilingual capabilities for universal information extraction. Firstly, it standardizes the representation of multilingual schemas using Python classes, ensuring a consistent ontology across different languages. Then, IE across languages is formulated as a unified code generation task. Secondly, we enhance the model’s cross-lingual transferability through IE cross-lingual alignment instruction tuning on a translated instance prediction task we proposed. During this phase, we also construct a high-quality and diverse bilingual IE parallel dataset with $\mathbf{257k}$ samples, called ParallelNER, synthesized by our proposed robust three-stage pipeline, with manual annotation to ensure quality. Although without training in $\mathbf{29}$ unseen languages, KnowCoder-X surpasses ChatGPT by $\mathbf{30.17\%}$ and SoTA by $\mathbf{20.03\%}$, thereby demonstrating superior cross-lingual IE capabilities. Comprehensive evaluations on $\mathbf{64}$ IE benchmarks in Chinese and English under various settings demonstrate that KnowCoder-X significantly enhances cross-lingual IE transfer through boosting the IE alignment. Our code and dataset are available at: https://github.com/ICT-GoKnow/KnowCoder.
@@
KnowCoder-X: Boosting Multilingual Information Extraction via Code
@@ 1 Introduction
Cross-lingual Information Extraction (IE) focuses on automatically extracting structured information from texts in unseen languages, following manually designed schemas. Cross-lingual schema understanding and representation have long been a challenge. Large Language Models (LLMs), trained on massive multilingual corpora, have significantly advanced multilingual language processing (Brown et al., 2020; Touvron et al., 2023a, b). Previous studies (Zhang et al., 2023; Qi et al., 2023; Wang et al., 2024; Gao et al., 2024; Li et al., 2024a) have shown that LLMs exhibit spontaneous cross-lingual alignment to facilitate the transfer of abilities and knowledge across languages. Our findings suggest the presence of this alignment in IE, indicating a strong potential for improving the IE cross-lingual transfer. We first define IE parallel data, which is across various languages that share the same schema, sentences, and extracted instances. Previous studies used label projection to generate such data to alleviate the challenges of IE in low-resource Kolluru et al. (2022); Hennig et al. (2023); Rios et al. (2024). We construct an IE parallel dataset in Chinese based on CoNLL 2003 Sang and Meulder (2003) manually, to evaluate the LLaMA2-7B trained on English Named Entity Recognition (NER) datasets used in KnowCoder Li et al. (2024b). Table 1 shows the results. After training on English IE datasets, there was a notable enhancement in the parallel dataset of Chinese IE ($5.9$$\rightarrow$52.752.752.752.7), strongly supporting the presence of IE cross-lingual alignment. However, the significant performance gaps between the two languages ($95.1$ vs $52.7$) indicate that the alignment remains weak. To address this, we propose two strategies to enhance IE cross-lingual alignment in LLMs by aligning the schema and extraction process across different languages, respectively:
Model CoNLL 2003 CoNLL 2003 (zh)
LLaMA2-7B $26.9$ $5.9$
LLaMA2-7B (w/ training) $95.1$ $52.7$
Table 1: Results on CoNLL 2003 and CoNLL 2003 (zh). The latter is the constructed Chinese parallel dataset.
Firstly, we standardize IE schemas across languages, particularly non-English ones, into unified Python classes. X-GEAR Huang et al. (2022) introduced a language-agnostic output template to maintain multilingual IE uniformity, aiding cross-lingual transfer. However, this uniformity is not thorough due to the absence of a schema in the input context, which hinders both cross-lingual transfer and multilingual schema understanding. Moreover, its highly customized input and output limit scalability to UIE and LLM-based IE. Thus, we introduce a unified UIE schema representation in the input and output, ensuring thorough uniformity in multilingual IE to improve cross-lingual alignment within LLMs. We also observe that object-oriented features in code are well-suited for unified schema representation and knowledge sharing across languages. Although code-based approaches Wang et al. (2022); Li et al. (2024b); Liu et al. (2025) have shown promising results in IE, their effectiveness in cross-lingual IE still needs to be validated. Therefore, we leverage Python classes to represent multilingual schemas, ensuring consistent ontology representation across languages and facilitating efficient IE knowledge transfer. The consistent schema allows the model to learn and efficiently share the knowledge of the same ontology across languages.
@@
Secondly, we introduce an IE cross-lingual alignment instruction tuning phase. We first propose a task that predicts target language instances, aiding in the alignment of the extraction process (see Figure 1 for an example). Using prediction tasks with parallel data to enhance multilingual alignment through instruction tuning is a widely adopted paradigm (Zhu et al., 2023; Gao et al., 2024). The key challenge is ensuring that extracted instances maintain semantic consistency with their source-language counterparts while appearing in the target language sentence. Unlike directly predicting complete IE parallel data, which focuses on aligning sentences, our method prioritizes the alignment of translated instances, which is the central goal of IE cross-lingual alignment.
@@
To use high-quality IE parallel data for the phase, we propose an automatic three-stage LLM-based pipeline, which achieves $99\%$ average faithfulness across $10$ languages evaluated on the label projection benchmark WikiANN Pan et al. (2017). Utilizing the pipeline, we construct a high-quality and diverse NER English-Chinese parallel dataset ParallelNER, using the GPT-4o series model supplemented with manual annotation. By distilling the IE alignment capabilities of advanced proprietary models, the IE cross-lingual alignment phase enhances cross-lingual generalization in IE tasks.
@@
Ultimately, we obtain the KnowCoder-X through Chinese and English IE instruction tuning. We evaluate the KnowCoder-X on $64$ benchmarks for the NER, Relation Extraction (RE), Event Detection (ED), and Event Argument Extraction (EAE) tasks in Chinese and English. KnowCoder-X exhibits superior cross-lingual generalization across $29$ unseen diverse languages, and surpasses ChatGPT by $30.17\%$ and SoTA by $20.03\%$. Moreover, it achieved an impressive average improvement of $11.43\%$ over the SoTA across $20$ low-resource African languages. In the supervised evaluation, KnowCoder-X consistently ranks within the top-$2$ results across $40$ (of $42$) benchmarks. Notably, KnowCoder-X achieves SoTA across all Chinese IE benchmarks, effectively demonstrating the strength of knowledge transfer from English IE through our method. Our contributions can be summarized as:
@@
  • The code-based multilingual IE method of KnowCoder-X unifies the representation of schemas across different languages, thereby boosting cross-lingual transfer.
@@
  • We incorporate an IE cross-lingual alignment phase to improve cross-lingual transfer. This phase involves instruction tuning on a newly proposed translated instance prediction task.
@@
  • We propose a robust three-stage IE parallel data construction pipeline and construct a high-quality bilingual parallel NER dataset ParallelNER with $257,190$ samples to provide valuable resources to the community.
@@
2 Related Work
Transfer Learning for cross-lingual IE   Huang et al. (2017) introduced a zero-shot learning method for EE, facilitating generalization across unseen languages. CLaP Huang et al. (2022) utilized generative language models for cross-lingual EAE. Prompt-XRE Hsu et al. (2023) enhanced cross-lingual RE through prompt-based learning, reducing reliance on large annotated datasets. Zubillaga et al. (2024) emphasized the significance of typology in cross-lingual EE, particularly for low-resource languages like Basque. In contrast, KnowCoder-X focuses on leveraging the spontaneous alignment of LLMs to facilitate cross-lingual IE transfer on all IE tasks.
LLMs for Multilingual IE   YAYI-UIE Xiao et al. (2023) proposes a chat-enhanced instruction tuning framework for UIE. IEPILE Gui et al. (2024) collects a comprehensive bilingual IE instruction corpus, and B2NER Yang et al. (2024) further designs a universal entity taxonomy. However, previous works overlook the mutual influence mechanisms between schemas in different languages and lack a unified representation.
Code-Based IE   Code-based IE aims to formulate the IE as a code generation task. CodeIE Li et al. (2023) first leverages the Code-LLM and recasts IE tasks into Python function completion tasks. Code4Struct Wang et al. (2022) represents the schema in Python classes for the EAE task. Code4UIE Guo et al. (2023) further constructs code-based schema for all the UIE tasks. Besides, KnowCoder Li et al. (2024b) utilizes class inheritance, class methods, and type hints to further enrich the schema representation. GoLLIE Sainz et al. (2024) incorporates guidelines within class comments. However, these works focus on the English IE task, and we are dedicated to exploring the significant role of code in multilingual IE.
@@ 3 Code-Based Multilingual IE
KnowCoder-X formulates the IE tasks across various languages using a unified code generation framework. In this section, we introduce the instruction and completion of the proposed code-based multilingual IE. The instruction consists of two parts (§3.1): (1) schema representation code across languages, including class definition and comments; (2) task prompts for completion. Subsequently, we present the completion format (§3.2).
@@ 3.1 Instruction Format
Class Definition   The IE schema comprises multiple concepts, each of which consists of a name and a set of attributes. The three fundamental concepts of IE schemas [Entity, Relation, Event] are first defined as base classes. For each concept of a schema, a corresponding class is subsequently defined. Each concept inherits from its respective base class by default. Then, we define the attributes of each concept (such as argument roles) as input arguments of the constructor __init__(). Specially, for IE in non-English languages, we first create a mapping from the original schema to the English schema. The mapping standardizes the representation of the same ontology across different languages by utilizing the unified Python class. For example, the entity concept {CJK}UTF8mj사람 in Korean and {CJK}UTF8gkai人物 in Chinese are both mapped to the class PER(Entity), ensuring consistent semantic representation across linguistic boundaries to align the schema in different languages.
Class Comments   We utilize class comments to provide clear definitions of concepts similar to GoLLIE Sainz et al. (2024) and KnowCoder Li et al. (2024b) for the understanding of concepts in different languages. Class comments comprise two components: 1) examples contain instances corresponding to the concept. 2) a description that explains the concept; For the examples part, we sample up to $10$ instances for each concept by frequency. For the description part, we follow the annotation guidelines of datasets by default. We propose an LLM-based approach to generate descriptions for datasets where guidelines are unavailable. To enable the LLM to better understand the concepts and the annotation style of the dataset, we first sample $10$ instances to prompt the LLM to summarize an initial description. To avoid providing excessive examples at once to hinder effective induction summarization, we then sample $20$ instances for the LLM to iteratively refine the description based on each one. Specifically, for each instance, if it cannot be categorized into the corresponding concept according to the description, the LLM will adjust the inappropriate parts of the current description accordingly. We use the GPT-4o-2024-08-06 as the LLM. Appendix A provides details on the description generation procedure.
Task Prompt   The task prompt includes a docstring defining the task and a string variable sentence that holds the text. The prompt is then concatenated with the schema to form the instruction.
@@ 3.2 Completion Format
The completion represents the final output generated by the model, which begins after the assignment operation (i.e., results =). Each completion is a list, where each element is an instantiated object, representing the extracted structured knowledge instance. Additionally, we utilize the exec() function in Python to execute the instructions and completions of all code-based data in this work to ensure quality. We also use this function to execute the generated code to extract the results.
4 IE Cross-Lingual Alignment
KnowCoder-X adopts a two-phase instruction tuning framework. The first phase is the IE cross-lingual alignment phase, where we propose the task and data. Subsequently, we conduct instruction tuning on Chinese and English UIE tasks with $46$ IE datasets to obtain KnowCoder-X. Appendix H shows examples of instruction tuning data. Appendix G shows the detailed dataset statistics.
In this section, we introduce the IE cross-lingual alignment phase. First, we introduce the instruction and completion of the translated spans prediction task (§4.1). To obtain high-quality IE parallel data for this training phase, we propose a three-stage automatic pipeline for constructing IE parallel data (§4.2) and creating ParallelNER (§4.3).
4.1 Translated Instances Prediction Task
@@
       // Obtain Prompt for Joint Translation.
@@
                  // Obtain Prompt for Span Rephrase. $i^{cor}_{m}=$ LLM ($prompt$) ;
@@
       $prompt=$ TemplateS ($I^{tgt},s^{src}$) ;
      // Obtain Prompt for Sentence Rephrase.
      $s^{tgt}=$ LLM ($prompt$); return $s^{tgt}$ ;
@@
Algorithm 1 Sample Translation
4.2 IE Parallel Data Construction Pipeline
The construction of IE parallel data involves two key components: sample translation (i.e., label projection) and schema translation. The sample consists of sentences and spans, where spans refer to the annotated segments within the sentence that are relevant to the IE task Kripke and Munitz (1971); Chen and Yuille (2004). Given the sentence $s^{src}$ and span list $I^{src}=\{i^{src}_{1},i^{src}_{2},...,i^{src}_{n}\}$ in the source language, label projection aims to obtain the sentence $s^{tgt}$ and span list $I^{tgt}=\{i^{tgt}_{1},i^{tgt}_{2},...,i^{tgt}_{n}\}$ in the target language. Schema translation can be carried out through machine translation and manual annotation. The primary challenge lies in the sample translation, which involves two key issues Parekh et al. (2024): 1) Inaccurate Translation: Either the spans or the sentences are inaccurately translated; 2) Missing Spans: The translated spans may be missing from the translated sentence.
Traditional label projection methods Dou and Neubig (2021); Chen et al. (2023); Parekh et al. (2024), such as CLaP, typically first generate $s^{tgt}$, and then obtain $I^{tgt}$ through label retrieval, word alignment, or contextual translation. However, these methods have a critical flaw: errors introduced during the generation of $s^{tgt}$ are propagated into the process of obtaining $I^{tgt}$. For example, during the translation of $s^{src}$ to get $s^{tgt}$, $i^{src}_{k}$ may be mistranslated or merged, especially when $i^{src}_{k}$ is a pronoun such as “his”, leading to the failure in obtaining the correct $i^{tgt}_{k}$. To address the issue, we propose a three-stage pipeline that implements joint translation and multi-grained rephrasing. Appendix I shows all prompts used in the pipeline.
Stage 1: Joint Translation   We propose the first stage to ensure semantic consistency between the translated sentence $s^{tgt}$ and the span list $I^{tgt}$. Compared to previous methods, where $s^{tgt}$ and $I^{tgt}$ are obtained sequentially, this stage utilizes LLMs to simultaneously translate and generate both $s^{tgt}$ and $I^{tgt}$, thereby reducing error accumulation through parallel processing. This stage provides $s^{src}$ and $I^{src}$ as a comprehensive context that introduces inner constraints on the two key issues, thus enabling a more accurate and natural translation. Specifically, we use a predefined template to format the sentence $s^{src}$ along with spans $I^{src}$ to serve as a prompt for the LLM to perform joint translation.